From: STOSC::SMM "On-Duty SMM x5074" 25-JAN-1993 20:34:52.31 To: @[SMM.DIS]HST_STAT CC: Subj: HST STATUS REPORT, 01/22 - 01/25 HUBBLE SPACE TELESCOPE DAILY REPORT #808 PERIOD COVERED: 0800 EST 01/22/93 - 0800 EST 01/25/93 Daily Status Report as of 025/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/UV2 1095 (Variability of High Luminosity Stars) The High Resolution Spectrograph was used to make ultraviolet observations of the vary massive high luminosity star HD15570. The observations were executed as scheduled, and no problems were reported. 1.2 Completed GHRS/2 3578 (Line Eclipse Mapping of Accretion Disk Wind) The High Resolution Spectrograph was used to make four spectral observations of the eclipse of the accretion disk of the cataclysmic variable star UX Ursa Majoris. All four observations were received and displayed. All of the spectral observations were reported to look normal. 1.3 Completed GHRS/2 4462 (SSA A2 Offset Verification) The High Resolution Spectrograph was used to make three positional observations of HD141556 to measure the position of the Small Science Aperture (SSA) with respect to the Large Science Aperture (LSA). All three to the target acquisitions and spectral observations were successful. 1.4 Completed PC 4347 (Flood Retention Test) The Planetary Camera was used to make three Earth flat calibration observations using filter F439W. These observations will be compared with those made in the past using the same filter, to determine how much of the last ultraviolet flood still remains. The observations were monitored in the OSS, and no problems were reported. 1.5 Completed Two Sets of GHRS/2 2537 (Physical Conditions of Interstellar Gas) The High Resolution Spectrograph was used to make high signal to noise ratio spectral observations of gas within 150 light years using star HD102647 (aka, Beta Leo, Denobola) to study the hot and cold regions of the interstellar medium. The spectral observations looked normal. The wave calibration images were contaminated by light from the target (which entered the HRS through the Small Science Aperture). 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed WFC-PC 4109 (Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make internal flat observations to check for different forms of contamination (i.e., daisies and measles). The OSS monitored to observations and assessed the data quality. No problems were reported. 1.8 Completed Two Sets of PC 3848 (Separate Light Curves of Pluto and Charon) The Planetary Camera was used to make two observations of the planet Pluto, and its moon Charon (each individually). Each of the images were taken using a comparison star which was about 25 arcseconds from the targets. The observations were completed as scheduled, and no problems were reported. 1.9 Completed FOS/RD 4052 (Inner Regions of Quasars) The Faint Object Spectrograph was used to make four ultraviolet observations of the inner regions of quasar B20923+39. The observations were received and displayed. They were reported to look a little noisy, but were otherwise okay. 1.10 Completed WFC 3917 (HST Medium Deep Survey, Part 1) The Wide Field Camera was used to make two high latitude observations which were taken in parallel with the prime science observations. Both observations were received and displayed in the OSS. They were reported to look okay. 1.11 Completed WFC 4216 (Calibration of Supernova) The Wide Field Camera was used to take three images of object IC4182. The images were received and displayed, and reported to look normal. 1.12 Completed WFC 2227 (Determine Extragalactic Distance Scale: M-81) The Wide Field Camera was used to make observations to the galaxy NGC-3031 (aka, M-81) to search for variable Cepheid stars using both filter F555W and F785LP individually. The observations were executed as scheduled, and no problems were reported. 1.13 Completed GHRS/2 3444 (Gas in Individual Interstellar Clouds) The High Resolution Spectrograph was used to make observations of early type star HD36486 to study the interstellar clouds which lie along the line of sight to that star. The observations were monitored in the OSS, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 15 Successful : 15 Scheduled Reacquisitions : 28 Successful : 28 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 25 Successful : 26 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 4152 (Observations to MOnitor Desorption) 3.2 Complete FOC/96 4069 (Far-UV Spectra of High Red-Shift Quasars) 3.3 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.4 Complete FOS/RD 3840 (Evolution of Carbon, Nitrogen and Oxygen in ST) 3.5 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.6 Complete WFC 4105 (HST Medium-Deep Survey: High Latitude) 3.7 Complete FOS/RD 4698 (FOS Dark Monitoring) 3.8 Complete GHRS/2 3444 (Gas in Interstellar Clouds) 3.9 Start of Zero-Gyro (On-Orbit) Test 026/1200Z through 028/0000Z. 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. The Zero-Gyro (On-Orbit) Test will be held from Tuesday 026/1200Z through Thursday 028/0000Z. The Fight Readiness Review meeting (FRR) was held today Friday Jan. 22 . -END-